A machine that fills bottles with a beverage has a fill volume whose mean is 22.01 ounces, with a standard deviation of 0.02 ounces. A case consists of 24 bottles randomly sampled from the output of the machine.

Find the mean of the total volume of the beverage in the case. Round the answer to two decimal places.

Find the standard deviation of the total volume of the beverage in the case.

Find of the mean of the average volume per bottle of the beverage in the case. Round the answer to two decimal places.

Find the standard deviation of the volume per bottle of the beverage in the case.

How many bottles must be included in a case for the standard deviation of the average volume per bottle to be 0.0025 ounces? Round the final answer to the next whole number.
Answer #1

1)

mean of the total volume of the beverage in the case =22.01*24=528.24

standard deviation of the total volume of the beverage in the case=0.02*√24=0.0980

2)

mean of the average volume per bottle of the beverage in the case =22.01

standard deviation of the volume per bottle of the beverage in the case =0.02/√24 =0.0041

3)

number of required bottles =(0.02/0.0025)^2=64